In a 4.0-kilometer race, a runner completes the first kilometer in 5.9 minutes, the second kilometer in 6.2 minutes, the third kilometer in 6.3 minutes, and the final kilometer in 6.0 minutes. the average speed of the runner for the race is approximately
1
Expert's answer
2020-10-08T18:13:36-0400

"v_{avr}=\frac{s_1+s_2+s_3+s_4 }{t_1+t_2+t_3+t_4}=\frac{1+1+1+1}{5.9+6.2.+6.3+6.0}=0.164(km/min)"